OpenStudy (anonymous):

integrate x^4e^(x/2)

OpenStudy (jhannybean):

\[\huge \int x^{{4e}^{x/2}}dx\]

OpenStudy (mimi_x3):

integration by parts pls

OpenStudy (mimi_x3):

wait

OpenStudy (anonymous):

integration by triple personality disorder

OpenStudy (jhannybean):

@Mimi_x3 go for it.

OpenStudy (anonymous):

you misspelled the second word honey

OpenStudy (jhannybean):

:) ...

OpenStudy (jhannybean):

@Michele_Laino Haha fun integral..

OpenStudy (mimi_x3):

could it be lel \[\int x^4 * e^{x/2}dx\]

OpenStudy (mimi_x3):

seens more plausible like this pls

OpenStudy (jhannybean):

How did you split it like that when it's a power of a power?

OpenStudy (mimi_x3):

the OP hasn't explicitly stated what the q is.

OpenStudy (jhannybean):

OH NEVERMIND. I wrote it wrong -.-

OpenStudy (mimi_x3):

integrate x^4e^(x/2) can be interpreated differently

OpenStudy (jhannybean):

It is infact \[\large \int x^4 e^{x/2}dx\]

OpenStudy (mimi_x3):

I WROTE THAT ALREEADY

OpenStudy (mimi_x3):

AND IBP PLS

OpenStudy (jhannybean):

Use LIPETs. \[u=x^4 ~,~ dv = e^{x/2}\]

OpenStudy (jhannybean):

Hehehe Mimi

OpenStudy (anonymous):

@Mimi_x3 Forgive me for letting the beast out of cage

ganeshie8 (ganeshie8):

\[\int x^ne^x dx = x^ne^x - n\int x^{n-1}e^x ~dx\]

OpenStudy (dan815):